At what rate of interest will a sum of money amount to Rs.5310 in 2 years and Rs.6120 in 4 years? Plz. give solution.
Answer by nerdybill(7384)

At what rate of interest will a sum of money amount to Rs.5310 in 2 years and Rs.6120 in 4 years? Plz. give solution.
.
Interest formula:
A = P(1+r)^t
.
Let P = initial sum of money
and r = rate of interest
then from "Rs.5310 in 2 years" we get
5310 = P(1+r)^2 (equation 1)
.
and from "Rs.6120 in 4 years" we get
6120 = P(1+r)^4 (equation 2)
.
solve equation 1 for P:
5310 = P(1+r)^2
5310/(1+r)^2 = P
.
substitute above into equation 2 and solve for r:
6120 = P(1+r)^4
6120 = 5310/(1+r)^2*(1+r)^4
6120/5310 = (1+r)^2
6120/5310 = (1+r)^2
1.1525423729 = (1+r)^2
1.0736 = 1+r
.0736 = r
or
7.4% = r